# Curbing PBHs with PTAs

### v1

@@ V.1 PTA observations
$W_{\rm D}(f)$ $\propto\delta\left(f\right)\,,$ (32)
$W_{\rm S}(f)$ $\propto{\rm sinc}^{2}\left(\pi\frac{f}{\Delta f}\right)\,,$ (33)
$W_{\rm TH}(f)$ $\propto\theta\left(f+\frac{\Delta f}{2}\right)\theta\left(\frac{\Delta f}{2}-f\right)\,,$ (34)
@@
which correspond to an infinitely accurate frequency resolution, a sharp binning of frequencies and a time-domain top-hat filter on $h_{\rm c}(t)$ with duration $T$, respectively. The window functions are normalized so that $\int{\rm d}\ln f\,W(f-f_{i})=1$. These window functions represent different limiting cases and can thus capture systematics related to the spectral resolution not otherwise accessible in our simplified analysis. As it can be seen from Fig. 1, the effect is expectedly larger at large scales and for heavy PBHs as these are mostly related to the signal at small frequencies. For instance, at $k\simeq 10^{6}\,{\rm{Mpc}}^{-1}$, $W_{\rm D}(f)$ gives a mildly stronger constraint on $A$ when compared to the other choices, as it can resolve sharp peaks in the SIGW spectrum. This is reflected in the constraints on the PBH abundance, as shown on the right panel of Fig. 1. This behaviour is reverted at even smaller $k_{*}$, and $W_{\rm D}(f)$ is less constraining than the other choices. For the remainder of the analysis, we will consider the Dirac window function $W_{\rm D}(f)$.
@@
Figure 2: Left panel: Constraints from NANOGrav15 Agazie et al. (2023a) on the amplitude of a log-normal curvature power spectrum, Eq. (1), with $\Delta=0.5$ (blue), and on a broken power-law curvature power spectrum, Eq. (2), with $\alpha=4$, $\gamma=1$ for $\beta=3$ (pink) and $\beta=0.5$ (red), assuming Gaussian fluctuations. The horizontal lines correspond to $f_{\rm PBH}=1$ using threshold statistics (solid) and theory of peaks (dashed), respectively, given the non-linearity-only scenario. Right panel: The implied NANOGrav15 constraints on PBH abundance.
@@
Figure 3: Left panels: Constraints from NANOGrav15 Agazie et al. (2023a) on the amplitude of a log-normal curvature power spectrum (1) with $\Delta=0.5$ (top panels) and on a broken power-law curvature power spectrum (2) with $\alpha=4$, $\gamma=1$ for $\beta=3$ (middle panels) and $\beta=0.5$ (bottom panels), assuming quadratic primordial NGs only with $|f_{\rm NL}|\in[0,10]$. The nearly horizontal lines correspond to $f_{\rm PBH}=1$ using threshold statistics with several $f_{NL}$ values. Right panels: Inferred constraints on the PBH abundance using threshold statistics.
@@
Another important theoretical uncertainty when estimating constraints on the PBH abundance arises due to the choice of the prescription when computing the PBH abundance. The right panels of Figs. 1 and 2 show the constraints on the PBH abundance arising when the abundance is computed using threshold statistics (solid lines) and peaks theory (dashed lines). In both figures, primordial NGs are neglected ($f_{\rm NL}=0$) so the NGs arise only from the non-linear relation between density and curvature perturbations. As one can see, almost the entire parameter space accessible by LVK via GWs from PBH binary mergers Hütsi et al. (2021); Romero-Rodriguez et al. (2022); Franciolini et al. (2022); Andrés-Carcasona et al. (2024) is excluded when threshold statistics is assumed. On the other hand, with the peaks theory, only a small portion of the parameter space relevant to LVK is constrained. Despite this significant theoretical uncertainty, PTA observations can constrain the PBH abundance in both cases.
@@
The dependence of the constraints on the shape of the power spectrum is illustrated in Fig. 2. While relatively narrow spectra, such as the LN and the BPL1 cases, produce similar results, broader spectra, represented by the BPL2 case are constrained also at larger scales due to the broadness of the spectrum. We remark that the constraints at large scales $k_{*}\lesssim 5\times 10^{6}{\rm Mpc}^{-1}$ arise mostly from the first bin by NANOGrav15. This is also the region in which the SIGW are most strongly constrained. In a similar vein, the constraints on the PBH abundance in the threshold statistics case are quite similar for all spectra except at large masses $\langle M_{\rm PBH}\rangle$. In particular, the strongest constraints touch the dashed region in Fig. 2 which corresponds to having less than a single PBH within the current Hubble volume. This line is given by $f_{\rm PBH}4\pi\Omega_{\rm DM}M_{\rm pl}^{2}/H_{0}<\langle M_{\rm PBH}\rangle$.
@@
As shown on the left panels of Fig. 3, for small $f_{\rm NL}$ we observe that the corrections to the SIGW affect only amplitudes at the tails, both at smaller and higher scales. A larger positive $f_{\rm NL}$, on the other hand, further shifts the constant $f_{\rm PBH}$ curves downward. As expected, this reduces the allowed amplitude $A$ for a given $k_{*}$. Conversely, for a fixed PBH abundance, positive primordial NGs decrease the required amplitude $A$ with respect to the Gaussian case. Since the latter effect is larger than the reduction of the SIGW strength, positive primordial NGs can completely remove constraints on the PBH abundance.
The situation differs when considering negative NGs. Since negative NGs imply an upward shift in the power spectrum amplitude in order to achieve the same PBH abundance, the constraints become more stringent.
@@
One might mistakenly assume that by increasing the negative value of the coefficient $f_{\rm NL}$, the required power spectral amplitude for the same PBH abundance would always rise, tightening the constraints indefinitely. However, this is not the case. As shown in Ref. Franciolini et al. (2023a), the required amplitude actually reaches a maximum for $f_{\rm NL}\simeq-2$ and then decreases, eventually approaching the amplitude of the Gaussian case for very large negative values of $f_{\rm NL}$. Consequently, the constraints for the case of $f_{\rm NL}=-2$ can be considered as the most stringent.
@@
Nevertheless, this claim should be taken cum grano salis. Indeed, it is important to note that the prescription outlined in Ref. Musco et al. (2021) to compute the threshold for PBH collapse only accounts for NGs arising from the non-linear relation between the density contrast and the curvature perturbations. In principle, also primordial NGs beyond the quadratic approximation should be taken into account when computing the threshold value. Following Refs. Kehagias et al. (2019); Escrivà et al. (2022), it appears that the effect on the threshold is small only for positive NGs and for small negative NGs ($f_{\rm NL}\gtrsim-2$) and at most of the order of a few percent. For more negative NGs, there are huge uncertainties on the exact value of the threshold, which could potentially modify the constraints introduced above [^5] One alternative approach is to use the averaged value over a sphere of radius equal to the location of the maximum of the compaction function, $\bar{C}_{\rm th}=2/5$, since this value is fully non-perturbative and independent of NGs Escrivà et al. (2020); Kehagias et al. (2024). As discussed in Ref. Ianniccari et al. (2024), this requires determining the statistics of the curvature perturbation and computing the connected cumulants, which is a complex task and is left for future work..
For completeness, we analyze a common scenario for PBH formation.
@@
In Ultra slow roll (USR) models of single-field inflation, the peak in $\mathcal{P}_{\zeta}$ arises from a brief phase of ultra-slow-roll followed by constant-roll inflation dual to it Atal and Germani (2019); Biagetti et al. (2018); Karam et al. (2023a). In this case, the NGs can be related to the large $k$ spectral slope Atal et al. (2019); Tomberg (2023),
$\zeta=-\frac{2}{\beta}\log\left(1-\frac{\beta}{2}\zeta_{\rm G}\right).$ (35)
In Fig. 4 we show a similar plot as in Fig. 2 but computing the abundance of PBHs using the full non-gaussian relation reported in Eq.(35) for the two broken power law cases reported in Tab. 1. We compute the abundance using exclusively the method based on the threshold statistics on the compaction function, and as we can see from the plot, with respect to the Gaussian approximation the constraints are highly relaxed and, differently from the Gaussian case, they are of the same order of magnitude. The reason is the following: expanding Eq. (35) at second order, we get $f_{\rm NL}=5\beta/12$, then using the reference points $\beta=3$ and $\beta=0.5$ shown in Fig. 4, one finds $f_{\rm NL}=5/4$ and $f_{\rm NL}=5/24$. As a consequence, the narrower power spectrum, which has stronger constraints in the Gaussian approximation, due to larger primordial non-gaussianities, has constraints similar to the broader case.
Moreover, with respect the impact on the PBH abundance, the NG corrections to the SIGW are irrelevant.
@@
Figure 4: The NANOGrav15 constraints on the PBH abundance using threshold statistics assuming the two cases for the broken power law reported in Tab. 1 assuming the gaussian approximation (solid) and with the full non-gaussian relation in Eq. (35).
@@
V.2 $\mu$ distortions and PBH seeds for SMBHs
At the largest scales, the primordial power spectrum is strongly constrained by the CMB observations Aghanim et al. (2020) and the Lyman-$\alpha$ forest data Bird et al. (2011). The CMB observations strongly constrain the curvature power spectrum at scales $10^{-4}\mathrm{Mpc}^{-1}\lesssim k\lesssim 1\mathrm{Mpc}^{-1}$. At redshifts $z\lesssim 10^{6}$, energy injections into the primordial plasma cause persisting spectral distortions in the CMB. These distortions are divided into chemical potential $\mu$-type distortions created at early times and Compton $y$-type distortions created at $z\lesssim 5\times 10^{4}$. For a given curvature power spectrum $\mathcal{P}_{\zeta}(k)$ the spectral distortions are Chluba et al. (2012b); Chluba and Grin (2013)
$X=\int_{-\infty}^{\infty}\frac{\mathrm{d}k}{k}\mathcal{P}_{\zeta}(k)W_{X}(k)$ (36)
with $X=\mu,y$ where $k_{\min}=1$ $\mathrm{Mpc}^{-1}$ and the window functions can be approximated by
$W_{\mu}(k)=2.2\left[e^{-\frac{(\hat{k}/1360)^{2}}{1+(\hat{k}/260)^{0.6}+\hat{k}/340}}-e^{-(\hat{k}/32)^{2}}\right],$ (37)
$W_{y}(k)=0.4e^{-(\hat{k}/32)^{2}}$ (38)
with $\hat{k}=k/\left(1\mathrm{Mpc}^{-1}\right)$. The COBE/Firas observations constrain the $\mu$ distortions as $\mu\leq 4.7\times 10^{-5}$ Bianchini and Fabbian (2022) and $y\leq 1.5\times 10^{-5}$ at $95\%$ confidence level Fixsen et al. (1996). The situation is summarised in Fig. 5. Including NGs modifies only the tails of the constraints and not the most constrained region Sharma et al. (2024).
@@ V.2 $\mu$ distortions and PBH seeds for SMBHs
We focus on the threshold statistics approach and we compute the mass fraction $\beta$ as a function of the parameters $f_{\rm NL}$ and $g_{\rm NL}$, assuming the power series ansatz as in Eq. 5. The power series expansion holds until we are in the perturbative regime, that is, the terms are ordered hierarchically with the higher orders being typically smaller than the lower ones. For a narrow power spectrum, this translates into a constraint for the amplitude of the power spectrum, $(3/5)\left|f_{\mathrm{NL}}\right|A^{1/2}\ll 1$, and $(9/25)\left|g_{\mathrm{NL}}\right|A\ll 1$.
@@
Indeed, the origin of SMBHs presents a significant challenge in the field of astrophysics. Although it is well known that they occupy the centres of most galaxies Kormendy and Richstone (1995); Magorrian et al. (1998); Richstone et al. (1998), the processes leading to their formation remain unclear.
In the context of PBHs, even a minimal presence of massive PBHs within the mass range $10^{3}-10^{6}\,M_{\odot}$ could potentially seed for SMBHs Duechting (2004); Kohri et al. (2014); Bernal et al. (2018). As a naive computation to estimate the necessary abundance of these seeds, we follow Refs. Vaskonen and Veermäe (2021); Serpico et al. (2020). Considering that SMBHs make up approximately $\mathcal{O}(10^{-4})$ of the stellar mass in their host galaxies Reines and Volonteri (2015), and that stars account for roughly $\mathcal{O}(10^{-2})$ of the total cosmic matter content Fukugita et al. (1998), we deduce that the overall SMBH density is about $10^{6}$ times less than the dark matter density.
@@
$f_{\mathrm{PBH}}\sim\mathcal{O}(10^{-6})\times\left\langle M_{\mathrm{PBH}}\right\rangle/M_{\rm SMBH}\,,$ (39)
with $\left\langle M_{\mathrm{PBH}}\right\rangle>10^{3}M_{\odot}$. Nevertheless, the scales related to the formation of SMBHs are strongly constrained by the analysis of the CMB spectral distortion by the FIRAS collaboration Fixsen et al. (1996); Chluba et al. (2012a, b); Chluba and Grin (2013); Bianchini and Fabbian (2022), and in the gaussian approximation, the corresponding abundance of PBHs is too small to furnish a primordial origin for Supermassive black holes seeds.
As shown in the left panel of Fig. 6, in agreement with the literature Ünal et al. (2021); Nakama et al. (2018); Hooper et al. (2024); Byrnes et al. (2024), quadratic primordial NGs are not enough to produce a sizeable amount of PBHs to be the seed of SMBHs. As a consequence, differently from Ref. Byrnes et al. (2

### v2

@@ 5.1 PTA observations
$W_{\rm D}(f)$ $\propto\delta\left(f\right)\,,$ (5.2)
$W_{\rm S}(f)$ $\propto{\rm sinc}^{2}\left(\pi\frac{f}{\Delta f}\right)\,,$ (5.3)
$W_{\rm TH}(f)$ $\propto\theta\left(f+\frac{\Delta f}{2}\right)\theta\left(\frac{\Delta f}{2}-f\right)\,,$ (5.4)
@@
which correspond to an infinitely accurate frequency resolution, a sharp binning of frequencies and a time-domain top-hat filter on $h_{\rm c}(t)$ with duration $T$, respectively. The window functions are normalized so that $\int{\rm d}\ln f\,W(f-f_{i})=1$. These window functions represent different limiting cases and can thus capture systematics related to the spectral resolution not otherwise accessible in our simplified analysis. As seen from Fig. 2, the effect is expectedly larger at large scales and for heavy PBHs as these are mostly related to the signal at small frequencies. For instance, at $k\simeq 10^{6}$ ${\rm{Mpc}}^{-1}$, $W_{\rm D}(f)$ gives a mildly stronger constraint on $A$ when compared to the other choices, as it can resolve sharp peaks in the SIGW spectrum. This is reflected in the constraints on the PBH abundance, as shown on the right panel of Fig. 2. This behaviour is reverted at even smaller $k_{*}$, and $W_{\rm D}(f)$ is less constraining than the other choices. For the remainder of the analysis, we will consider the Dirac window function $W_{\rm D}(f)$.
@@
Figure 3: Left panel: Constraints from NANOGrav15 [29] on the amplitude of a log-normal curvature power spectrum, Eq. (2.1), with $\Delta=0.5$ (blue), and on a broken power-law curvature power spectrum, Eq. (2.2), with $\alpha=4$, $\gamma=1$ for $\beta=3$ (pink) and $\beta=0.5$ (red), assuming Gaussian fluctuations. The horizontal lines correspond to $f_{\rm PBH}=1$ using threshold statistics (solid) and theory of peaks (dashed), respectively, given the non-linearity-only scenario. Right panel: The implied NANOGrav15 constraints on PBH abundance.
@@
Figure 4: Left panels: Constraints from NANOGrav15 [29] on the amplitude of a log-normal curvature power spectrum (2.1) with $\Delta=0.5$ (top panels) and on a broken power-law curvature power spectrum (2.2) with $\alpha=4$, $\gamma=1$ for $\beta=3$ (middle panels) and $\beta=0.5$ (bottom panels), assuming quadratic primordial NGs only with $|f_{\rm NL}|\in[0,10]$. The nearly horizontal lines correspond to $f_{\rm PBH}=1$ using threshold statistics with several $f_{NL}$ values. Right panels: Inferred constraints on the PBH abundance using threshold statistics.
@@
Another important theoretical uncertainty when estimating constraints on the PBH abundance arises due to the choice of the prescription when computing the PBH abundance. The right panels of Figs. 2 and 3 show the constraints on the PBH abundance arising when the abundance is computed using threshold statistics (solid lines) and peaks theory (dashed lines). In both figures, primordial NGs are neglected ($f_{\rm NL}=0$) so the NGs arise only from the non-linear relation between density and curvature perturbations. As one can see, almost the entire parameter space accessible by LVK via GWs from PBH binary mergers [161, 162, 137, 123] is excluded when threshold statistics is assumed. On the other hand, with the peaks theory, only a small portion of the parameter space relevant to LVK is constrained. Despite this significant theoretical uncertainty, PTA observations can constrain the PBH abundance in both cases.
@@
The dependence of the constraints on the shape of the power spectrum is illustrated in Fig. 3. While relatively narrow spectra, such as the LN and the BPL1 cases, produce similar results, broader spectra, represented by the BPL2 case, are constrained also at larger scales due to the broadness of the spectrum. We remark that the constraints at large scales $k_{*}\lesssim 5\times 10^{6}{\rm Mpc}^{-1}$ arise mostly from the first bin by NANOGrav15. This is also the region in which the SIGW are most strongly constrained. In a similar vein, the constraints on the PBH abundance in the threshold statistics case are quite similar for all spectra except at large masses $\langle M_{\rm PBH}\rangle$. In particular, the strongest constraints touch the dashed region in Fig. 3 which corresponds to having less than a single PBH within the current Hubble volume. This line is given by $f_{\rm PBH}4\pi\Omega_{\rm DM}M_{\rm pl}^{2}/H_{0}<\langle M_{\rm PBH}\rangle$.
@@
As shown on the left panels of Fig. 4, for small $f_{\rm NL}$ we observe that the corrections to the SIGW affect only amplitudes at the tails, both at smaller and higher scales. Larger NGs ($f_{\rm NL}\geq 5$), on the other hand, shift the entire SIGW constraint curves. The effect on the $f_{\rm PBH}$ constraints, shown in the right panels of Fig. 4, can be qualitatively understood by comparing the shift of $f_{\rm PBH}=1$ curves when compared to the SIGW constraint: Since the effect on $f_{\rm PBH}=1$ is larger than the strengthening of the SIGW constraints, positive primordial NGs can completely remove constraints on the PBH abundance. We find that for $f_{\rm NL}=5$, the PTA constraints on $f_{\rm PBH}$ are eliminated in the wide BPL2 case and become weaker than the constraints from other observables in the narrow LN and BPL1 cases.
The opposite is observed with negative NGs – since they strengthen the constraints on SIGW in a similar way to positive NGs, but require a higher amplitude $A$ to achieve the same PBH abundance, the constraints on $f_{\rm PBH}$ become more stringent. This is indicated by the green line in Fig. 4. In particular, even a mild negative NG ($f_{\rm NL}=-2$) could exclude the existence of any PBHs in our present Hubble volume in the mass range $2-100M_{\odot}$, given a narrow spectrum as in the LN and BPL1 cases.
@@
However, one might mistakenly assume that by increasing the negative value of the coefficient $f_{\rm NL}$, the required power spectral amplitude for the same PBH abundance would always rise, tightening the constraints indefinitely. However, this is not the case, as the required amplitude reaches a maximum at $f_{\rm NL}\simeq-2$ and then decreases, eventually approaching the amplitude of the Gaussian case for very large negative values of $f_{\rm NL}$ [46]. Consequently, the constraints for the case of $f_{\rm NL}=-2$ can be considered the most stringent.
@@
Nevertheless, this claim should be taken cum grano salis. Indeed, it is important to note that the prescription outlined in Ref. [121] to compute the threshold for PBH collapse only accounts for NGs arising from the non-linear relation between the density contrast and the curvature perturbations. In principle, also primordial NGs beyond the quadratic approximation should be taken into account when computing the threshold value. Following Refs. [163, 164], it appears that their effect on the threshold is small when the NGs are positive or mildly negative ($f_{\rm NL}\gtrsim-2$). In such cases, the threshold receives corrections of at most a few percent. For stronger negative NGs, the uncertainties on the exact value of the threshold can be sizeable, and could thus potentially modify the constraints introduced above [^6] One alternative approach is to use the averaged value over a sphere of radius equal to the location of the maximum of the compaction function, $\bar{C}_{\rm th}=2/5$, since this value is fully non-perturbative and independent of NGs [165, 166]. As discussed in Ref. [125], this requires determining the statistics of the curvature perturbation and computing the connected cumulants, which is a complex task and is left for future work..
Although evaluating the NG corrections to the formation threshold is beyond the scope of this work, we can estimate the impact on the PBH abundance due to a few percent correction on $\mathcal{C}_{\rm th}$. In detail, we re-computed the constraints on $f_{\rm PBH}$ considering a 5% change in $\mathcal{C}_{\rm th}$ for the three benchmark models in Table 1. In all the cases, the constraints on the abundance were affected by at most two orders of magnitude. Therefore, in the light of Fig. 4, we expect that perturbative corrections on $\mathcal{C}_{\rm th}$ would not induce significant changes to the PTA constraints on $f_{\rm PBH}$ and thus our main conclusions would not be strongly affected.
@@
For completeness, we analyze a specific but common scenario for PBH formation. In Ultra slow roll (USR) models of single-field inflation, the peak in $\mathcal{P}_{\zeta}$ arises from a brief phase of ultra-slow-roll which is typically followed by slow-roll or constant-roll inflation dual to it [167, 168, 57]. In this case, the NGs can be related to the large $k$ spectral slope generated during the last inflationary phase [169, 170],
$\zeta=-\frac{2}{\beta}\log\left(1-\frac{\beta}{2}\zeta_{\rm G}\right).$ (5.5)
In Fig. 5, we show the abundance of PBHs using the full NG relation Eq. (5.5) for the two broken power law cases reported in Tab. 1. The PBH abundance is computed using threshold statistics and compared with the Gaussian approximation. As seen from Fig. 5, NGs can significantly affect the constraints on $f_{\rm PBH}$ when the spectrum is narrow (in the BPL1 case). For wider spectra (BPL2 case), the NG-induced suppression is milder and thus the final constraint is stronger. To make contact with the $f_{\rm NL}$ description of NGs, expanding Eq. (5.5) yields $f_{\rm NL}=5\beta/12$, so that $\beta=3$ and $\beta=0.5$ correspond to $f_{\rm NL}=5/4$ and $f_{\rm NL}=5/24$, thus the NGs are relatively small and the NG corrections to SIGW have a subdominant effect. Consequently, for USR models, PTAs are more effective in constraining PBHs from wide curvature power spectra due to the weaker NG-induced suppression.
@@
Figure 5: The NANOGrav15 constraints on the PBH abundance using threshold statistics assuming the two cases for the broken power law reported in Tab. 1 assuming the Gaussian approximation (solid) and with the full NG relation in Eq. (5.5).
@@
5.2 $\mu$ distortions and PBH seeds for SMBHs
At the largest scales, the primordial power spectrum is strongly constrained by the CMB observations [152] and the Lyman-$\alpha$ forest data [181] [^7] For recent works on how future CMB experiments can probe the range of scales related to PTA see Refs. [182, 183]. The CMB observations strongly constrain the curvature power spectrum at scales $10^{-4}\mathrm{Mpc}^{-1}\lesssim k\lesssim 1\mathrm{Mpc}^{-1}$. At redshifts $z\lesssim 10^{6}$, energy injections into the primordial plasma cause persisting spectral distortions in the CMB. These distortions are divided into chemical potential $\mu$-type distortions created at early times and Compton $y$-type distortions created at $z\lesssim 5\times 10^{4}$. For a given curvature power spectrum $\mathcal{P}_{\zeta}(k)$ the spectral distortions are [53, 54]
$X=\int_{k_{\rm min}}^{\infty}\frac{\mathrm{d}k}{k}\mathcal{P}_{\zeta}(k)W_{X}(k)$ (5.6)
with $X=\mu,y$, while $k_{\rm min}=1$ Mpc-1 and the window functions can be approximated by
$W_{\mu}(k)=2.2\left[e^{-\frac{(\hat{k}/1360)^{2}}{1+(\hat{k}/260)^{0.6}+\hat{k}/340}}-e^{-(\hat{k}/32)^{2}}\right],$ (5.7)
$W_{y}(k)=0.4e^{-(\hat{k}/32)^{2}}$ (5.8)
with $\hat{k}=k/\left(1\mathrm{Mpc}^{-1}\right)$. The COBE/Firas observations constrain the $\mu$ distortions as $\mu\leq 4.7\times 10^{-5}$ [55] and $y\leq 1.5\times 10^{-5}$ at $95\%$ confidence level [51]. The situation is summarised in Fig. 6. Including NGs modifies only the tails of the constraints and not the most constrained region [184].
@@ 5.2 $\mu$ distortions and PBH seeds for SMBHs
We focus on the threshold statistics approach and we compute the mass fraction $\beta$ as a function of the parameters $f_{\rm NL}$ and $g_{\rm NL}$, assuming the power series ansatz as in Eq. 2.5. The power series expansion holds until we are in the perturbative regime, that is, the terms are ordered hierarchically with the higher orders being typically smaller than the lower ones. For a narrow power spectrum, this translates into a constraint for the amplitude of the power spectrum, $(3/5)\left|f_{\mathrm{NL}}\right|A^{1/2}\ll 1$, and $(9/25)\left|g_{\mathrm{NL}}\right|A\ll 1$.
@@
Indeed, the origin of SMBHs presents a significant challenge in the field of astrophysics. Although it is well known that they occupy the centres of most galaxies [185, 186, 187], the processes leading to their formation remain unclear.
In the context of PBHs, even a minimal presence of massive PBHs within the mass range $10^{3}-10^{6}\,M_{\odot}$ could potentially seed for SMBHs [48, 188, 49]. As a naive computation to estimate the necessary abundance of these seeds, we follow Refs. [42, 50]. Considering that SMBHs make up approximately $\mathcal{O}(10^{-4})$ of the stellar mass in their host galaxies [189], and that stars account for roughly $\mathcal{O}(10^{-2})$ of the total cosmic matter content [190], we deduce that the overall SMBH density is about $10^{6}$ times less than the dark matter density.
@@
$f_{\mathrm{PBH}}\sim\mathcal{O}(10^{-6})\times\left\langle M_{\mathrm{PBH}}\right\rangle/M_{\rm SMBH}\,,$ (5.9)
with $\left\langle M_{\mathrm{PBH}}\right\rangle>10^{3}M_{\odot}$. Nevertheless, the scales related to the formation of SMBHs are strongly constrained by the analysis of the CMB spectral distortion by the FIRAS collaboration [51, 52, 53, 54, 55], and in the gaussian approximation, the corresponding abundance of PBHs is too small to furnish a primordial origin for Supermassive black holes seeds.
As shown in the left panel of Fig. 7, in agreement with the literature [191, 192, 193, 194], quadratic primordial NGs are not enough to produce a sizeable amount of PBHs to be the seed of SMBHs. Consequently, differently from Ref. [194], we find that already a sizeable cubic parameter $g_{\rm NL}$ is enough to overcome this hurdle. This discrepancy arises because their analysis is subject to a few simplifications: using the curvature perturbation instead of the density contrast to determine the collapse and neglecting the threshold’s dependence on the shape of the curvature power spectrum.
According to Fig. 7, assuming a typical SMBH mass of approximately $10^{8}$ $M_{\odot}$, condition (5.9), without the violation of the perturbative criterium, holds across the range of PBH masses analyzed.
6 Conclusions and outlook
In this work, we have updated the constraints on the amplitude of the curvature power spectrum, ensuring that the amplitude of the produced SIGWs does not exceed the signal recently detected by the NANOGrav collaboration.
When PBHs are formed via the collapse of sizeable curvature perturbations, the constraints on SIGW will infer an upper bound on the PBH abundance. We derived these bounds in various scenarios and by estimating PBH abundance using threshold statistics as well as the theory of peaks.
We found that, when using threshold statistics, only a small fraction of dark matter in the form of PBHs is available in the solar mass range regardless of the specific shape of the power spectrum. For narrow curvature power spectra, significant constraints $f_{\rm PBH}\lesssim 10^{-5}$ arise in the mass range $0.1-100M_{\rm\odot}$, while for broader spectra, such constraints arise for heavier $5-500M_{\rm\odot}$ PBHs.
These potential constraints can, however, be significantly relaxed in the presence of positive NGs, and are eliminated when $f_{\rm NL}\gtrsim 10$. However, we find that stellar mass PBHs tend to remain constrained when considering NGs present in typical USR models for PBH production. In these cases, wider spectra tend to infer strong constraints over a wider PBH mass range as they are associated with weaker NGs.
When the PBH abundance is estimated using the theory of peaks, however, constraints on $f_{\rm PBH}$ are strongly relaxed. Although they are not completely eliminated (in the absence of non-Gaussianities) they tend to be in the same order as existing constraints on the PBH abundance.
Finally, we discussed how large primordial NGs ($g_{\rm NL}\simeq 10^{5}$) permit the production of PBH seeds for SMBHs without violating perturbativity. Despite this possibility, the literature currently lacks explicit models capable of producing sufficiently large NGs. Indeed, in common scenarios, such as the curvaton and USR models, the general impact of the NGs is curtailed when one goes beyond the perturbative approach [63, 195].
To improve future estimates, it is crucial to refine the computation of the PBH abundance, which is still subject to a series of theoretical uncertainties.
Acknowledgments
We thank G. Franciolini and X. Pritchard for useful discussions and A. Ianniccari, D. Perrone and V. Vaskonen for comments on the draft. A.J.I. acknowledges the financial support provided under the “Progetti per Avvio alla Ricerca Tipo 1”, protocol number AR1231886850F568. H.V. is supported by the Estonian Research Council grants PSG869 and RVTT7 and the Center of Excellence program TK202.